Circular No. 8737

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 P1 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery observation

tabulated below) on CCD images obtained with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey, noting the object to have a faint 20′′-diameter coma in moonlight. Following posting on the ‘NEO Confirmation Page’, C. Jacques and E. Pimentel report that CCD images obtained at Belo Horizonte, Brazil, with a 0.30-m f/3 Schmidt-Cassegrain reflector on Aug. 8.0 UT show a coma diameter of 16′′.

2006 UT  α_{2000}  δ_{2000}  Mag.
Aug. 7.50872  16^{h}38^{m}58.^{s}70  −18^{o}05′46′′.3  17.3

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-P31.

T = 2007 June 17.713 TT  ω = 109.^{o}726
Ω = 268.367 }2000.0
q = 1.55508 AU  i = 143.691

COMET C/1998 J5 (SOHO) Further to IAUC 8735, the following Kreutz-sungrazing archival SOHO

comet was diffuse, elongated, and very faint.

Comet 1998 UT α_{2000} δ_{2000} Inst. F MPEC
C/1998 J5 May 12.227 3^{h}17.^{m}5 +16^{o}46′ C2 HS 2006-L20

COMET P/2006 HR_{30} (SIDING SPRING) Further to IAUC 8735, J. Licandro, Isaac Newton Group and Insti-

tuto de Astrofisica de Canarias; and N. Pinilla and M. Pedani, Centro Galileo Galilei, report that comet P/2006 HR_{30} shows no evidence of a coma on R-band images obtained on July 26.10 UT with the 3.6-m Telescopio Nazionale Galileo at La Palma; six co-added 60-s images show the comet as point-like (FWHM 1′′; ellipticity of 0.06, nearly equal to that of the surrounding stars). D. T. Durig (Sewanee, TN, 0.30-m reflector) writes that images obtained on Aug. 4.20–4.28 show a slight asymmetry or tail ∼ 10′′ long in p.a. 5^{o}–10^{o}, and P. C. Sherrod (Conway, AR, 0.41-m reflector) notes that the comet showed some nebulosity toward the north on images taken on Aug. 4.4.

2006 August 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green